# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v3

@@ 4.2 Style Transfer Results
Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$. This strong boost in robustness comes at a small price of base style capture fidelity. As observed in Figures 4 and 5, we can also treat subsampling as an introduction of noise. Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid. However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the biggest outlier in the dataset.
4.3 Quantitative Evaluation
User Study
To evaluate the utility of our approach under different DP and subsampling configurations, we conducted a user study with 25 participants. Each participant was shown reference images from the target dataset and asked to compare pairs of generated images, selecting the one that better captured the style of the reference images. Images were generated using 10 prompts and adapted TI embeddings for the @eveismyname and Paris 2024 Pictogram datasets, resulting in 20 groups of images. Each participant evaluated two groups, one randomly selected from each dataset, with comparisons focusing on model configurations differing by DP noise and subsampling size.
Survey results, summarized in Table 1 in Appendix A, align with our design goals. Participants showed no clear preference between regular TI and DPAgg-TI, suggesting that our privacy-preserving approach maintains perceptual quality. As expected, both DP noise and reduced subsampling size degraded style fidelity, consistent with the trade-offs inherent in differential privacy. Preferences at $\varepsilon=1$ were split, but subsampling was generally favored, reinforcing its role in reducing noise impact while preserving style.
Kernel Inception Distance
The Kernel Inception Distance (KID) (Bińkowski et al. 2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space. To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al. 2015), and their distributional differences are estimated. Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al. 2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al. 2024), making it more suitable for smaller datasets, as in our case.
We report KID scores for different parameters in Tables  2 and 3 (see Appendix B), showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy. Further discussion of these results is also provided in Appendix B.
4.4 Ablation Study: Textual Inversion with DP-SGD
A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training. We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets. We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset. In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 6. We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images. Additional results can be found in Appendix F.
5 Copyright Protection Implications
Our proposed mechanism can also be interpreted through the lens of copyright protection. This connection is grounded in the framework of Near Access-Freeness (NAF) (Vyas, Kakade, and Barak 2023), which evaluates whether a model’s outputs reveal undue influence from specific data points by comparing them to those from a safe model trained without access to the same data.
Since DPAgg-TI satisfies $\varepsilon$-DP, it also satisfies $\varepsilon$-NAF, which means the adapted model behaves similarly to one that never saw the private images, under the NAF criterion. However, we emphasize that this guarantee holds only within the NAF framework; it does not constitute a general claim about content similarity or legal compliance. Crucially, DPAgg-TI is designed to adapt to the style of private images, not their specific content. Prior work and legal precedent suggest that style imitation is generally considered fair use and does not constitute infringement (Vyas, Kakade, and Barak 2023). Thus, our mechanism aligns with the intended protections of NAF: it avoids memorization while still enabling meaningful personalization and stylistic adaptation. We defer the details of copyright protection to Appendix D.
6 Conclusion
Acknowledgements
We sincerely thank @eveismyname for providing her artwork for use in this study. We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study. Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
Ethical Statement
Appendix A User Study
@@
Table 1: Survey Results.
Figure 7: Samples of image sets used in our user study. Participants are asked to compare 2 images at a time.
A.1 Study Design and Objective
The user study aimed to assess the utility of our approach under different DP and subsampling configurations by evaluating the models’ ability to adapt to novel styles. The study involved 25 participants, each of whom was tasked with comparing images generated using various configurations and selecting the one that better captured the style of reference images.
A.2 Experimental Setup
Participants were shown reference images from two datasets:
  • The @eveismyname dataset of private artwork.
  • The Paris 2024 Pictogram dataset.
For each dataset, 10 prompts were used to generate images, resulting in 20 groups of images (10 prompts per dataset). Each group included images generated using the same prompt and dataset but with different model configurations. Configurations varied in the addition of DP noise and the size of subsampling.
  • Original Textual Inversion (TI)
  • DPAgg-TI ($\varepsilon=\infty$, no DP) w/o subsampling
  • DPAgg-TI ($\varepsilon=1$) without subsampling
  • No Adaptation
  • DPAgg-TI ($\varepsilon=\infty$, no DP) with subsampling $(m=8)$
  • DPAgg-TI ($\varepsilon=1$) with subsampling $(m=8)$
  • Style Guidance (SG)
A.3 Survey Procedure
Participants were asked to evaluate two groups of images: one randomly selected from the @eveismyname dataset and one from the Paris 2024 Pictogram dataset. For each group:
  1. 1.
    Participants were shown reference images from the target dataset.
  2. 2.
    They were presented with pairs of images generated using different model configurations for the same prompt.
  3. 3.
    Participants selected the image they felt better captured the style of the reference images.
A.4 Evaluation Metrics
The study focused on assessing:
  • Participants’ preference between regular TI and DPAgg-TI for style adaptation.
  • The impact of DP noise and subsampling size on the perceived utility of style transfer.
A.5 Results and Analysis
The results are summarized in Table 1. Key observations include:
  • Participants showed no clear preference between regular TI and DPAgg-TI in capturing styles for either dataset.
  • Both DP noise and reduced subsampling size decreased the perceived quality of style transfer.
  • Preferences were split between configurations with $\varepsilon=1$ with and without subsampling, though subsampling generally had favorable outcomes.
These findings highlight the trade-off between increased DP robustness and reduced utility, suggesting that the optimal configuration may depend on subjective preferences and specific application requirements.
Appendix B Kernel Inception Distance
@@
Table 2: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl). Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
@@
Table 3: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl). Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
Our results indicate that DPAgg-TI preserves the style transfer fidelity of TI while also ensuring differential privacy. Notably, for @eveismyname ($m=4$) at low privacy budgets, we observe even lower KID values than standard TI, suggesting enhanced style alignment. Similarly, results for the Paris 2024 dataset follow a comparable trend, with DPAgg-TI achieving KID scores similar to TI at low privacy budgets. However, the overall KID scores for this dataset remain high within the context of diffusion model style transfer.
Upon inspecting the generated images (Figure 8), we hypothesize that the abstract and out-of-distribution nature of the Paris 2024 images poses a challenge for the Inception network, leading to less meaningful feature embeddings. This likely inflates the measured embedding distances between generated and reference images, resulting in higher-than-expected KID values.
For KID evaluations, we used prompts similar to those employed during TI training: “A painting/icon in the style of $S^{*}$”. Consistent with the training image captions, these prompts do not specify a subject. For each parameter configuration, we generate 100 images and compute KID by repeatedly subsampling the larger of the real and generated sets to match the size of the smaller set, 100 times, then averaging the resulting KID scores.
@@
Appendix C Differentially Private Adaptation via Style Guidance
C.1 Background: Denoising Diffusion Implicit Models
@@
Denoising Diffusion Implicit Models (DDIM) sampling (Song, Meng, and Ermon 2021) uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$. The reverse diffusion process for DDIM sampling can be formalized as follows:
@@
C.2 Implementation
We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al. 2024). Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
@@ C.2 Implementation
Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder. The forward guidance process is defined as follows:
@@
C.3 Style Transfer Results
@@
The reduced effectiveness of SG for style transfer may stem from its sensitivity to hyperparameters such as the guidance weight $w$, leading to instability. Although Bansal et al. (2024) proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application. Additionally, the CLIP embeddings may not retain enough stylistic detail after the aggregation.
C.4 Ablation
To better understand the limited effectiveness of style guidance in our experiments, despite its success in (Bansal et al. 2024), we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection(Innat 2022) (Figure 10). Unlike the @eveismyname and Paris 2024 datasets, it is highly likely that Stable Diffusion has been trained on these images. Additionally,  Bansal et al. (2024) demonstrated successful adaptation towards the style of Van Gogh’s Starry Night as a single reference image, making this dataset a reasonable interpolation between their successful results and our more limited findings.

### v4

@@ IV-B Style Transfer Results
Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to with respect to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$. This strong boost in robustness comes at a small price of base style capture fidelity. As observed in Figures 4 and 5, we can also treat subsampling as an introduction of noise. Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid. However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the biggest outlier in the dataset.
Figure 6: Samples of image sets used in our user study. Participants were asked to compare 2 images at a time.
TABLE I: Survey Results.
@@
IV-C User Study
To evaluate our approach under different DP and subsampling configurations, we conducted a user study with 25 participants. The goal was to assess whether DPAgg-TI preserves perceptual quality while offering privacy guarantees.
IV-C1 Study Design and Setup
Participants were shown reference images from two datasets: the @eveismyname dataset of private artwork and the Paris 2024 Pictogram dataset. For each dataset, we used 10 prompts to generate images, resulting in 20 groups in total. Each group included images produced under different configurations, including regular TI, DPAgg-TI with and without DP noise, with and without subsampling ($m=8$), style guidance (see Appendix A), and a no-adaptation baseline.
IV-C2 Survey Procedure
Each participant evaluated two groups, one randomly selected from each dataset. For each group, participants were first shown the reference images, then asked to compare pairs of generated images produced with different configurations (see Figure 6). For each pair, they indicated which image better captured the reference style, or marked the choice as “unsure.”
IV-C3 Results and Analysis
Survey results are summarized in Table I. Participants showed no clear preference between regular TI and DPAgg-TI, suggesting that our privacy-preserving approach maintains perceptual quality. As expected, both DP noise and smaller subsampling size degraded style fidelity, consistent with the trade-offs inherent in differential privacy. At $\varepsilon=1$, preferences were split between configurations with and without subsampling, although the subsampling variant was generally favored.
Overall, the findings highlight that DPAgg-TI achieves perceptual quality comparable to regular TI, while subsampling serves as an effective mechanism to balance privacy and stylistic fidelity.
IV-D Kernel Inception Distance
The Kernel Inception Distance (KID) [26] is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space. To compute KID, images generated by the model and real training images are passed through an Inception network [27], and their distributional differences are estimated. Unlike the more commonly used Fréchet Inception Distance (FID) [28], KID is an unbiased estimator of the true divergence between the learned and target distributions [29], making it more suitable for smaller datasets, as in our case.
TABLE II: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl). Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
@@ IV-D Kernel Inception Distance
TABLE III: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl). Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
@@
We report KID scores for different parameters in Tables  II and III. Our results indicate that DPAgg-TI preserves the style transfer fidelity of TI while also ensuring differential privacy. Notably, for @eveismyname ($m=4$) at low privacy budgets, we observe even lower KID values than standard TI, suggesting enhanced style alignment. Similarly, results for the Paris 2024 dataset follow a comparable trend, with DPAgg-TI achieving KID scores similar to TI at low privacy budgets. However, the overall KID scores for this dataset remain high within the context of diffusion model style transfer.
Upon inspecting the generated images (Figure 7), we hypothesize that the abstract and out-of-distribution nature of the Paris 2024 images poses a challenge for the Inception network, leading to less meaningful feature embeddings. This likely inflates the measured embedding distances between generated and reference images, resulting in unusually high KID values.
For KID evaluations, we used prompts similar to those employed during TI training: “A painting/icon in the style of $S^{*}$”. Consistent with the training image captions, these prompts do not specify a subject. For each parameter configuration, we generate 100 images and compute KID by repeatedly subsampling the larger of the real and generated sets to match the size of the smaller set 100 times, then averaging the resulting KID scores.
@@
IV-E Ablation Study: Textual Inversion with DP-SGD
A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training. We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets. We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset. In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 8. We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images. Additional results can be found in Appendix C.
V Discussion
V-A Copyright Protection Implications
Our proposed mechanism can also be interpreted through the lens of copyright protection. This connection is grounded in the framework of Near Access-Freeness (NAF) [6], which evaluates whether a model’s outputs reveal undue influence from specific data points by comparing them to those from a safe model trained without access to the same data.
Modern generative models typically produce outputs via randomized sampling. Leveraging this inherent randomness, Vyas et al. [6] introduced NAF as a metric to quantify the similarity between a model’s output and copyrighted content. The key idea is to compare the output distribution of a potentially infringing model to that of a safe model—one trained without access to the target content. A canonical example is the leave-one-out-safe model, trained on the full dataset excluding $x$. Since $\mathsf{safe}(x)$ lacks access to $x$, the probability that it generates content resembling $x$ is expected to be small; any such resemblance is considered fortuitous.
Definition 2 (Near Access-Freeness [6]).
Let $\mathcal{C}$ be a set of copyrighted samples and $\mathcal{W}$ a set of generative models. Given a mapping $\mathsf{safe}:\mathcal{C}\rightarrow\mathcal{W}$ and a divergence measure $\Delta$, we say a model $w\in\mathcal{W}$ is $k_{y}$-near access-free (or $k_{y}$-NAF) on prompt $y\in\mathcal{Y}$ if for every $x\in\mathcal{C}$,
$\Delta\!\left(P_{w}(\cdot|y)\,\|\,P_{\mathsf{safe}(x)}(\cdot|y)\right)\leq k_{y}.$
NAF is closely related to concepts in DP. Depending on the divergence measure $\Delta$, NAF resembles different DP variants—for example, $\varepsilon$-DP when $\Delta=\Delta_{\mathrm{max}}$ [10], and $(1,\varepsilon)$-Rényi DP when $\Delta=\Delta_{\mathrm{KL}}$. Translating DP to generative models yields this definition:
Definition 3 (Differentially Private Generation (DPG)).
Let $S$ and $S^{\prime}$ be neighboring datasets. Denote by $P_{S}(\cdot|y)$ the distribution over outputs generated by a model trained on, or adapted from, $S$ with algorithm $\mathcal{A}$, where randomness includes both training and generation. The generation satisfies $\varepsilon$-Differentially Private Generation ($\varepsilon$-DPG) if for every $y\in\mathcal{Y}$,
Here, neighboring datasets differ by a single data point (or privacy unit). If the training process is $\varepsilon$-DP, then the outputs naturally satisfy $\varepsilon$-DPG via the data processing inequality. One benefit of DPG is the flexibility to add noise during generation rather than training, potentially improving the utility–privacy tradeoff. However, there are notable distinctions: $\varepsilon$-DP offers protection under arbitrary post-processing and multiple outputs, whereas $\varepsilon$-DPG only guarantees privacy for single outputs. Also, under DP, the trained model can be released, but under DPG, only the outputs are safe to share. Elkin-Koren et al. [30] highlight further differences: NAF is one-sided—comparing a model to a fixed safe reference—whereas DPG is symmetric. This asymmetry in NAF can enable better utility. Additionally, NAF allows more flexibility in choosing the safe model, which can be exploited in algorithm design.
Since DPAgg-TI satisfies $\varepsilon$-DP, it also satisfies $\varepsilon$-NAF under the leave-one-out-safe model. Within the NAF framework, this means the adapted model behaves similarly to one that never saw the private images. Importantly, this guarantee is meaningful only within NAF; it does not imply broader legal immunity or empirical indistinguishability from the original content. However, it allows us to argue that any close resemblance between outputs and private training data is no more likely than would be expected from a model with no access to that data.
Finally, the goal of DPAgg-TI is to adapt to the style of private image sets, not their precise content. This distinction matters: pure style imitation (without reproducing protectable expression or “substantive elements”) is often argued to be non-infringing in many creative contexts, though the legal status is jurisdiction- and fact-dependent; particularly in artistic and creative contexts. As discussed in Elkin-Koren et al. [30] and legal analyses such as Carlini et al. [5], generating new content in the style of a work, without reproducing its substantive elements, is generally not considered copyright infringement. Therefore, the use of DPAgg-TI to learn and reproduce stylistic attributes does not contradict the spirit or intent of the NAF framework. Instead, it offers a promising direction for responsibly fine-tuning generative models on private or copyrighted sources while respecting both privacy and intellectual property boundaries.
Remark 1 (Scope of Protection and Artist-Level Extension).
We provide record-level DP: it limits leakage or reconstruction of any individual private image, not an artist’s entire style. Consequently, it yields a corresponding NAF-style guarantee at the per-image level (under the chosen safe reference). This interpretation should be understood strictly within NAF and does not constitute a general copyright compliance claim. The DP guarantee continues to apply under targeted prompts: conditioning on detailed descriptions can increase the likelihood of reproducing a specific private work by at most an $e^{\varepsilon}$ factor (up to $\delta$). Artist-level (user-level) DP is conceptually possible by treating each artist as one unit and privately aggregating artist-level embeddings (e.g., via a DP mean mechanism), but typically requires stronger noise and may reduce utility; we leave a full exploration to future work.
V-B Limitations
DPAgg-TI is designed for the low-data, strong-privacy regime, where the number of private images is small ($n\approx 100$) and per-record protection with $\varepsilon<5$ matters. For large datasets with moderate subsampling, DP-SGD on the full model may become more efficient and could provide better utility. We explicitly position our method for scenarios where DP-SGD is known to struggle: strong privacy guarantees with limited training data.
We acknowledge that in moderate privacy regimes ($5\leq\varepsilon\leq 10$) with larger batch sizes and careful tuning, DP-SGD with parameter-efficient fine-tuning methods might perform better than our approach. However, in our experiments, applying DP-SGD to regular TI even at $\varepsilon=20$ with carefully tuned hyperparameters (learning rate, batch size, scheduler) still required prohibitively high noise levels to satisfy the privacy accountant, preventing meaningful learning. Moreover, DP-SGD typically involves multiple training iterations, so the effective noise further accumulates due to composition across epochs, making convergence extremely difficult in the strong-privacy regime ($\varepsilon<5$) with approximately 100 images.
VI Conclusion
Ethical Statement
Acknowledgments
We sincerely thank @eveismyname for providing her artwork for use in this study. We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study. Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google. This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2025-23525649).
Appendix A Differentially Private Adaptation via Style Guidance
A-A Background: Denoising Diffusion Implicit Models
@@
Denoising Diffusion Implicit Models (DDIM) sampling [15] uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$. The reverse diffusion process for DDIM sampling can be formalized as
@@
A-B Implementation
We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance [31]. Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
@@ A-B Implementation
Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder. The forward guidance process is defined as
@@
A-C Style Transfer Results
@@
The reduced effectiveness of SG for style transfer may stem from its sensitivity to hyperparameters such as the guidance weight $w$, leading to instability. Although Bansal et al. [31] proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application. Additionally, the CLIP embeddings may not retain enough stylistic detail after the aggregation.
A-D Ablation
To better understand the limited effectiveness of style guidance in our experiments, despite its success in Bansal et al. [31], we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection[32] (Figure 10). Unlike the @eveismyname and Paris 2024 datasets, it is highly likely that Stable Diffusion has been trained on these images. Additionally, Bansal et al. [31] demonstrated successful adaptation towards the style of Van Gogh’s Starry Night as a single reference image, making this dataset a reasonable interpolation between their successful results and our more limited findings.